Word problem. "The area of a certain square is 5 more than its perimeter. What is the length of each side of the square." and if i could see the what the equation looks like to, that would be wonderful. Thanks!
Answer by Theo(3458)

area of a square is given by the equation:

A = s^2

where s is one side of the square, and A is the area.

The perimeter of a square is given by the equation:

P = 4*s

where s is one side of the square, and P is the perimeter.

If the area of the square is 5 more than the perimeter, this means that:

A = P + 5

which means that:

s^2 = 4*s + 5

subtract 4*s + 5 from both sides of the equation to get:

s^2 - 4*s - 5 = 0

this factors out to be:

(s-5) * (s+1) = 0

this makes:

s = 5 and s = -1

s is the side of the square and can't be negative, so s must equal 5.

each side of the square is equal to 5

the area of the square is 5 * 5 = 25

the perimeter of the square is 4 * 5 = 20

25 - 20 = 5 so the area of the square is 5 more than the perimeter of the square.